# DT-UFC: Universal Large Model Feature Coding via Peaky-to-Balanced Distribution Transformation

### v1

@@ Abstract.
Like image coding in visual data transmission, feature coding is essential for the distributed deployment of large models by significantly reducing transmission and storage overhead. However, prior studies have mostly targeted task- or model-specific scenarios, leaving the challenge of universal feature coding across diverse large models largely unaddressed. In this paper, we present the first systematic study on universal feature coding for large models. The key challenge lies in the inherently diverse and distributionally incompatible nature of features extracted from different models. For example, features from DINOv2 exhibit highly peaky, concentrated distributions, while those from Stable Diffusion 3 (SD3) are more dispersed and uniform. This distributional heterogeneity severely hampers both compression efficiency and cross-model generalization. To address this, we propose a learned peaky-to-balanced distribution transformation, which reshapes highly skewed feature distributions into a common, balanced target space. This transformation is non-uniform, data-driven, and plug-and-play, enabling effective alignment of heterogeneous distributions without modifying downstream codecs. With this alignment, a universal codec trained on the balanced target distribution can effectively generalize to features from different models and tasks. We validate our approach on three representative large models—LLaMA3, DINOv2, and SD3—across multiple tasks and modalities. Extensive experiments show that our method achieves notable improvements in both compression efficiency and cross-model generalization over task-specific baselines. All source code will be released for future research.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Dublin, Irelandisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Image compressionccs: Computing methodologies Image segmentationccs: Computing methodologies Natural language generation
Figure 1. The overall framework of DT-UFC. Large models generate heterogeneous features (red CDF curves). The proposed adaptive transformation module (blue curves) reshapes these features into a unified target distribution space. During inference, the decoded features are inverse-transformed and passed to task-specific heads to complete predictions. (Sec. 4.1)
@@ 1. Introduction
In recent years, large foundation models such as GTP-4 (Achiam et al., 2023), DeepSeek (Guo et al., 2025), and Gemini (Team et al., 2024) have demonstrated remarkable capabilities across a wide range of reasoning, discriminative, and generative tasks. As these models are increasingly integrated into distributed and resource-constrained systems (Tian et al., 2022; Sagawa* et al., 2020; Ye et al., 2024), the transmission of intermediate representations has become a bottleneck for scalable deployment. Similar to image and video coding, feature coding aims to encode intermediate features compactly while preserving semantic fidelity. However, existing feature coding approaches are predominantly model- or task-specific (Henzel et al., 2022; Wang et al., 2024; Ma et al., 2024; Alvar and Bajić, 2019; Yan et al., 2021), where the encoder is tailored to a particular feature distribution. Such specialization severely limits their universality in realistic scenarios where multiple large models coexist and are applied to various downstream tasks.
@@
Building on this insight, we propose DT-UFC, a novel framework for Universal Feature Coding via Peaky-to-Balanced Distribution Transformation. The core of DT-UFC is the learnable non-uniform peaky-to-balanced feature transformation that adaptively transforms diverse input feature distributions into a common, balanced target distribution. This transformation is designed to be plug-and-play and does not require any modification to the downstream codec. Once transformed, the features can be compressed by a universal codec and then inversely transformed for downstream tasks. This not only bridges the distribution gap between features but also enhances the expressiveness of the latent representation and improves entropy coding efficiency. To evaluate the generalization ability of DT-UFC, we conduct experiments across three representative tasks: common sense reasoning (CSR), semantic segmentation (Seg), and text-to-image generation (TTI). These tasks cover both discriminative and generative paradigms and span vision and language modalities, providing a comprehensive evaluation for universal feature coding. Our main contributions are summarized as follows:
@@ 2.1. Feature Coding
Feature coding, as a key branch of coding for machines, was originally introduced to compress intermediate features in collaborative intelligence scenarios, where deep models are split across edge and cloud. Early work in (Chen et al., 2019) evaluated coding performance using standard image codecs but suffered from low efficiency due to distribution mismatch. Subsequent research focused on task-specific feature coding (Choi and Bajić, 2018; Cai et al., 2022; Henzel et al., 2022; Wang et al., 2024; Ma et al., 2024; Gao et al., 2024b), where compression pipelines are tailored to single-task scenarios like image classification or detection. While effective, these methods lack generalizability across diverse tasks or models.
To broaden applicability, recent efforts explored multi-task feature coding (Alvar and Bajić, 2019; Yan et al., 2021; Alvar and Bajić, 2020; Feng et al., 2022; Gao et al., 2024a; Wang et al., 2022b; Zhang et al., 2021b; Guo et al., 2023; Wang et al., 2023a). For instance, (Yan et al., 2021) proposed a scalable coding scheme for coarse-to-fine tasks. However, these methods typically rely on task-specific decoders or side information and remain constrained to fixed sets of tasks or models. Recognizing the growing importance of feature coding, standardization efforts have also emerged, notably through MPEG’s ongoing work on feature coding standards (WG2, 2023).
Both academic research and standardization have predominantly focused on features extracted from CNN-based models (Zhang et al., 2021b; Kim et al., 2023; Misra et al., 2022; Liu et al., 2021). In contrast, the current landscape of artificial intelligence has been dominated by large models (Guo et al., 2025; Achiam et al., 2023; Bai et al., 2023; Oquab et al., 2023; Touvron et al., 2023; Esser et al., 2024). To bridge this gap, (Gao et al., 2024c) introduced the first large model feature coding benchmark, providing a dataset, unified test conditions, and baseline methods. This work highlights the critical role of feature coding in large model deployment. However, the proposed baselines require model-specific training and do not address the universal feature coding problem.
@@ 2.2. Feature Alignment
Feature alignment is a common technique in representation learning to reduce domain shift and improve generalization. It has been widely applied in domain adaptation (Lee and Lee, 2023; Yeh et al., 2021; Long et al., 2015; Sun and Saenko, 2016; Wang et al., 2023b), multi-modal learning (Radford et al., 2021; Li et al., 2021, 2022, 2023; Ge et al., 2021), and multi-task learning (Wang et al., 2022a; Senushkin et al., 2023), enabling models to share or transfer representations across distributions, modalities, or tasks.
@@
3. Challenges in Universal Feature Coding
3.1. Feature Distribution Analysis
@@ 3.2. Challenges in Universality
To validate this, we characterize the similarity between different feature distributions using Kullback-Leibler (KL) divergence. Each feature distribution is approximated using a discretized histogram. The KL Divergence is computed as:
(1) $D_{\mathrm{KL}}(p\|q)=\sum_{x}p\log\frac{p}{q}$
As shown in Table 1, original features from different models exhibit large KL divergence values, confirming their inherent distributional misalignment. This implies that an encoder trained on one feature distribution must allocate additional bitrates to accurately compress features from another distribution. In contrast, the KL divergence among transformed feature distributions is significantly lower, demonstrating that our proposed distribution transformation (see Section 4) effectively aligns statistical properties across different models. This alignment serves as a critical enabler for universal feature coding.
@@ 3.3. Challenges in Coding Performance
Beyond generalization, peaky input distributions also negatively impact the efficiency and expressiveness of the learned encoders. Specifically, they restrict the effective exploitation of the latent space, destabilize quantization, and reduce entropy coding efficiency.
@@ 4.2. Non-uniform Feature Transformation
  • The range of $\mathbf{f}$ is partitioned into $256$ non-overlapping intervals =$\mathcal{R}=\{R_{0},R_{1},\dots,R_{255}\}$, with each $R_{k}\subset\mathbb{R}$.
@@ 5.4. Cross-Task Universality
Baseline with Transformed Features (Sec. 5.5) 2.08 94.00 0.04 31.60 0.12 161.72
@@
DT-UFC with Baseline Features (Sec. 5.6) 0.30 6.60 0.05 37.32 0.08 166.57
@@ 5.6. Codec’s Robustness to Imperfect Inputs
Moreover, the smooth decline of metrics such as MSE and FID across bitrate indicates that codecs trained with transformed features have not merely memorized specific input patterns but have learned to model underlying feature structures more generally. Therefore, DT-UFC is not only effective under ideal conditions but also robust in real-world deployment scenarios where features may be distorted or misaligned. This resilience is critical for practical adoption.
@@ 5.7. Architecture-Agnostic Generalization
Table 6 shows that DT-UFC achieves nearly lossless classification accuracy and decreasing MSE across bitrates. This demonstrates the codec’s strong generalization from Transformer-derived to CNN-derived features. Importantly, this result highlights our method’s ability to align feature distributions across architectural paradigms, enabling universal feature compression across diverse model backbones. In addition, the generalization ability also indicates that the proposed transformation can align feature distributions across model architectures, enabling universal feature coding across Transformer-based and CNN-based backbones. The strong performance without retraining further validates the scalability and practicality of DT-UFC as a truly architecture-agnostic feature codec.

### v2

@@ Abstract.
Like image coding in visual data transmission, feature coding is essential for the distributed deployment of large models by significantly reducing transmission and storage burden. However, prior studies have mostly targeted task- or model-specific scenarios, leaving the challenge of universal feature coding across diverse large models largely unexplored. In this paper, we present the first systematic study on universal feature coding for large models. The key challenge lies in the inherently diverse and distributionally incompatible nature of features extracted from different models. For example, features from DINOv2 exhibit highly peaky, concentrated distributions, while those from Stable Diffusion 3 (SD3) are more dispersed and uniform. This distributional heterogeneity severely hampers both compression efficiency and cross-model generalization. To address this, we propose a learned peaky-to-balanced distribution transformation, which reshapes highly skewed feature distributions into a common, balanced target space. This transformation is non-uniform, data-driven, and plug-and-play, enabling effective alignment of heterogeneous distributions without modifying downstream codecs. With this alignment, a universal codec trained on the balanced target distribution can effectively generalize to features from different models and tasks. We validate our approach on three representative large models (LLaMA3, DINOv2, and SD3) across multiple tasks and modalities. Extensive experiments show that our method achieves notable improvements in both compression efficiency and cross-model generalization over task-specific baselines. All source code has been made available at https://github.com/chansongoal/DT-UFC.
@@
journalyear: 2025copyright: ccconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia (MM ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3755814isbn: 979-8-4007-2035-2/2025/10ccs: Computing methodologies Image compressionccs: Computing methodologies Computer visionccs: Computing methodologies Natural language processing
Figure 1. The overall framework of DT-UFC. Large models generate heterogeneous features (red CDF curves). The proposed adaptive transforms (blue curves) reshape these features into a unified target distribution space. The decoded features are inverse-transformed and passed to task-specific heads to complete predictions. (Sec. 4.1)
@@ 1. Introduction
In recent years, large foundation models such as GTP-4 (Achiam et al., 2023), DeepSeek (Guo et al., 2025), and Gemini (Team et al., 2024) have demonstrated remarkable capabilities across a wide range of reasoning, discriminative, and generative tasks. As these models are increasingly integrated into distributed and resource-constrained systems (Tian et al., 2022; Sagawa* et al., 2020; Ye et al., 2024), the transmission of intermediate representations has become a bottleneck for scalable deployment. Similar to image and video coding (Li et al., 2024c; Gao et al., 2022; Zuo et al., 2024; Mao et al., 2024; Li et al., 2024d; Tang et al., 2024; Li et al., 2024a, 2022a, b), feature coding aims to encode intermediate features compactly while preserving semantic fidelity. However, existing feature coding approaches are predominantly model- or task-specific (Henzel et al., 2022; Wang et al., 2024; Ma et al., 2024; Alvar and Bajić, 2019; Yan et al., 2021; Chen et al., 2024), where the encoder is tailored to a particular feature distribution. Such specialization severely limits their universality in realistic scenarios where multiple large models coexist and are applied to various downstream tasks.
@@
Building on this insight, we propose DT-UFC, a novel framework for Universal Feature Coding via Peaky-to-Balanced Distribution Transformation. The core of DT-UFC is the learnable non-uniform peaky-to-balanced transform that can adaptively transforms diverse input feature distributions into a common, balanced target distribution. This transformation is designed to be plug-and-play and does not require any modification to the downstream codec. Once transformed, the features can be compressed by a universal codec and then inversely transformed for downstream tasks. This not only bridges the distribution gap between features but also enhances the expressiveness of the latent representation and improves entropy coding efficiency. To evaluate the generalization ability of DT-UFC, we conduct experiments across three representative tasks: common sense reasoning (CSR), semantic segmentation (Seg), and text-to-image generation (TTI). These tasks cover both discriminative and generative paradigms and span vision and language modalities, providing a comprehensive evaluation for universal feature coding. Our main contributions are summarized as follows:
@@ 2.1. Feature Coding
Feature coding, as a key branch of coding for machines (Gao et al., 2023; Zhu et al., 2024; Yin et al., 2025; Shen et al., 2024a, b; Liu et al., 2024; Chen et al., 2023; Kao et al., 2024; Tian et al., 2023, 2024a, 2024b), was originally introduced to compress intermediate features in collaborative intelligence scenarios, where deep models are split across edge and cloud. Early work in (Chen et al., 2019) evaluated coding performance using standard image codecs but suffered from low efficiency due to distribution mismatch. Subsequent research focused on task-specific feature coding (Choi and Bajić, 2018; Cai et al., 2022; Henzel et al., 2022; Wang et al., 2024; Ma et al., 2024; Gao et al., 2024b), where compression pipelines are tailored to single-task scenarios like image classification or detection. While effective, these methods lack generalizability across diverse tasks or models.
To broaden applicability, recent efforts explored multi-task feature coding (Alvar and Bajić, 2019; Yan et al., 2021; Alvar and Bajić, 2020; Feng et al., 2022; Gao et al., 2024a, 2025; Wang et al., 2022b; Zhang et al., 2021b; Guo et al., 2023; Wang et al., 2023a). For instance, (Yan et al., 2021) proposed a scalable coding scheme for coarse-to-fine tasks. However, these methods typically rely on task-specific decoders or side information and remain constrained to fixed sets of tasks or models. Recognizing the growing importance of feature coding, standardization efforts have also emerged, notably through MPEG’s ongoing work on feature coding standards (WG2, 2023).
Both academic research and standardization have predominantly focused on features extracted from CNN-based models (Zhang et al., 2021b; Kim et al., 2023; Misra et al., 2022; Liu et al., 2021; Wu et al., 2025). In contrast, the current landscape of artificial intelligence has been dominated by large models (Guo et al., 2025; Achiam et al., 2023; Bai et al., 2023; Oquab et al., 2023; Touvron et al., 2023; Esser et al., 2024). To bridge this gap, (Gao et al., 2024c) introduced the first large model feature coding benchmark, providing a dataset, unified test conditions, and baseline methods. This work highlights the critical role of feature coding in large model deployment. However, the proposed baselines require model-specific training and do not address the universal feature coding problem.
@@ 2.2. Feature Alignment
Feature alignment is a common technique in representation learning to reduce domain shift and improve generalization. It has been widely applied in domain adaptation (Lee and Lee, 2023; Yeh et al., 2021; Long et al., 2015; Sun and Saenko, 2016; Wang et al., 2023b), multi-modal learning (Radford et al., 2021; Li et al., 2021, 2022b, 2023; Ge et al., 2021), and multi-task learning (Wang et al., 2022a; Senushkin et al., 2023), enabling models to share or transfer representations across distributions, modalities, or tasks.
@@
3. Challenges in Universal Feature Coding
3.1. Feature Distribution Analysis
@@ 3.2. Challenges in Universality
To validate this, we characterize the similarity between different feature distributions using Kullback-Leibler (KL) divergence. Each feature distribution is approximated using a discretized histogram. The KL divergence is computed as $D_{\mathrm{KL}}(p\|q)=\sum p\log(p/q)$. As shown in Table 1, original features from different models exhibit large KL divergence values, confirming their inherent distributional misalignment. This implies that an encoder trained on one feature distribution must allocate additional bitrates to accurately compress features from another distribution. In contrast, the KL divergence among transformed feature distributions is significantly lower, demonstrating that our proposed distribution transformation (see Section 4) effectively aligns statistical properties across different models. This alignment serves as a critical enabler for universal feature coding.
@@ 3.3. Challenges in Coding Performance
Beyond generalization, peaky input distributions also negatively impact the efficiency and expressiveness of learned encoders. Specifically, they restrict the effective exploitation of the latent space, destabilize quantization, and reduce entropy coding efficiency.
@@ 4.2. Non-uniform Feature Transformation
  • The range of $\mathbf{f}$ is partitioned into $256$ non-overlapping intervals $\mathcal{R}=\{R_{0},R_{1},\dots,R_{255}\}$, with each $R_{k}\subset\mathbb{R}$.
@@ 5.4. Cross-Task Universality
Baseline with
Transformed
Features
(Sec. 5.5)
2.08 94.00 0.04 31.60 0.12 161.72
@@
DT-UFC with
Baseline
Features
(Sec. 5.6)
0.30 6.60 0.05 37.32 0.08 166.57
@@ 5.6. Codec’s Robustness to Imperfect Inputs
Moreover, the decline of metrics such as MSE and FID across bitrate indicates that codecs trained with transformed features have not merely memorized specific input patterns but have learned to model underlying feature structures more generally. Therefore, DT-UFC is not only effective under ideal conditions but also robust in real-world deployment scenarios where features may be distorted or misaligned. This resilience is critical for practical adoption.
@@ 5.7. Architecture-Agnostic Generalization
Table 6 shows that DT-UFC achieves nearly lossless classification accuracy and decreasing MSE across bitrates. This demonstrates the codec’s strong generalization from Transformer features to CNN features. Importantly, this result highlights our method’s ability to align feature distributions across architectural paradigms, enabling universal feature compression across diverse model backbones. In addition, the generalization ability also indicates that the proposed transformation can align feature distributions across model architectures, enabling universal feature coding across Transformer-based and CNN-based backbones. The strong performance without retraining further validates the scalability and practicality of DT-UFC as a truly architecture-agnostic feature codec.
@@ 6. Conclusion
Acknowledgements.
This work was supported by the Ministry of Education of Singapore under Grant T2EP20123-0006. We acknowledge the support of GPU cluster built by MCC Lab of Information Science and Technology Institution, USTC.